OpenStudy (anonymous):

Solve for x: 2 over 5x + 2 = 10

OpenStudy (bahrom7893):

2/(5x+2) = 10?

OpenStudy (sgadi):

2=50x+20 x=-18/50=-0.36

OpenStudy (anonymous):

it is either 20 24 4 or 8

OpenStudy (sgadi):

I am sorry, answer is 3.6 the approximate values is 4. The answer I have given is by solving 2/(5x+2) = 10. If the answer is not matched, the question may not be correct.

OpenStudy (anonymous):

2 ----x + 2 = 10 this is the question 5

OpenStudy (sgadi):

(2/5)*x+2 = 10 =>2x/5 = 8 => 2x = 40 => x = 20

OpenStudy (anonymous):

\[\frac{2x}{5}+2=10\\ \frac{2x}{5}=8\\ 2x=8*5=40\\ x=20\]